# Dockformer: A transformer-based molecular docking paradigm for large-scale virtual screening

### v2

@@ IV-A Prediction Performance on Benchmark Datasets
In this section, the docking accuracy of Dockformer was evaluated on the PDBbind core set and PoseBusters dataset and compared with that of five commonly used optimization-based docking algorithms, namely, GOLD [6], Glide [32], LeDock [33], AutoDock Vina [34] and Mini Vina, and nine state-of-the-art DL-based docking approaches, i.e., DeepDock [20], EquiBind [13], TankBind [21], DiffDock [15], Uni-Mol [14], KarmaDock [22], GAABind [23], CarsiDock [24] and Umol [35].
@@
DL-based docking methods are susceptible to the input training sample distrubtion, and changing the search space dramatically decreases the model capacity. Therefore, the trained models provided by the official repositories with the default search spaces were used in the comparison study. Specifically, since EquiBind, DiffDock, TankBind and Umol are trained for blind docking tasks, their search spaces cover the entire crystal protein. The binding pockets of Uni-Mol, GAABind and Dockformer are defined as the protein residues within the range of from any heavy atom of a crystal ligand, whereas those of KarmaDock and CarsiDock are considered the protein residues within the range of and , respectively. DeepDock considers the protein surface mesh nodes within of any crystal ligand atom as inputs. In addition, the number of binding pocket boxes was set to for all conventional docking methods. The root mean square deviation (RMSD), which measures the geometric similarity between the predicted binding conformations and the crystal structures of the ligands, was used to evaluate the docking approaches. Generally, binding pose predictions are considered successful when their RMSDs are below the threshold of [36].
@@
As illustrated in Fig. 5(a), Dockformer achieves the highest docking success rates of 90.53% and 82.71% on the PDBbind core set and PoseBusters dataset, respectively. These rates are higher than those of all the baselines. CarsiDock is the second-best model, with slight success rate decreases of 1.76% and 7.01% on the two benchmarks, respectively. TankBind, EquiBind and DiffDock perform relatively poor because they were originally trained to solve blind docking tasks. In addition, the docking performance on the PoseBusters dataset is worse for all approaches but especially for DL-based methods. For example, KarmaDock achieves a success rate of 83.86% using the PDBbind core set and 46.73% using PoseBusters. This finding implies that some DL-based methods may not generalize well to unseen data because PoseBusters contains only complexes released since 2021. However, the accuracy of Dockformer ranks first on this dataset, suggesting its strong generalizability. Interestingly, most optimization-based docking methods achieve satisfactory and robust performance on both datasets. These methods are superior to those of most DL-based methods. LeDock achieves the second-best accuracy and is slightly inferior to Dockformer using PoseBusters. As a more comprehensive and effective method, the cumulative frequencies of the binding poses with the RMSDs from their corresponding crystal ligands for all the docking methods are plotted in Fig. 5(b). We find that the success rates of binding poses generated by GOLD and LeDock are higher than those of most DL-based docking approaches. However, Dockformer still performs very competitively with different RMSD cutoffs on both benchmarks, indicating its obvious superiority in terms of accuracy.
IV-B Computational Complexity versus Accuracy
@@
In addition to accuracy, computational complexity is an important performance indicator that requires attention, especially when the screening compound library becomes extremely large. Most docking methods cannot traverse the entire library within an acceptable running time. As depicted in Fig. 6, Dockformer yields the highest docking accuracy and requires the least amount of time using the PoseBusters dataset among all the docking methods. Although Dockformer has the largest DL network, end-to-end binding conformation generation results in low computational complexity in the inference process. Both LeDock and CarsiDock can also achieve competitive docking accuracies but with a long inference time. Thus, these approaches are unsuitable for LSVS tasks. Moreover, most DL-based docking approaches are more efficient than optimization-based approaches are but sacrifice gains in accuracy. In addition, the docking accuracy of DL-based methods improves as the size of the network architecture increases.
@@
The computational cost of optimization-based docking methods is substantially greater than that of DL-based methods. The adopted stochastic optimization algorithms require more computing resources to determine the position, orientation, and torsion angles of each ligand conformation according to the scoring functions. The DL-based methods, including EquiBind, TankBind, Uni-Mol, GAABind and CarsiDock, use DL algorithms to construct intra- and intermolecular distance maps and then adopt optimization algorithms to calculate the coordinates of ligand atoms for binding pose generation. These methods cannot abandon independent optimization procedures for each ligand, which are faster than those of optimization-based docking methods but still time-consuming because of the iterative gradient descent processes. Only Dockformer and KarmaDock use end-to-end network architecture modules to generate binding conformations in a batch fashion. These modules have prominent advantages in docking efficiency. A modified version of Dockformer termed Dockformer Raw was also evaluated in this experiment. Dockformer Raw uses the same geometry optimization strategy as that used in TankBind to predict binding conformations.
@@
The experimental results demonstrate that Dockformer Raw performs worse in terms of both accuracy and inference time. This finding highlights that the superiority of end-to-end structures can effectively mitigate the computational burden of iterative optimization. Notably, AlphaFold3 achieves a success rate of 90.2% on the PoseBusters dataset, a rate much higher than that achieved of each the docking methods mentioned in our experiments. However, even in the case of the least number of tokens, the inference time of AlphaFold3 is 22 on 16 A100 graphics processing units (GPUs). This time is much longer than that of most competitors because the model framework of AlphaFold3 is much larger [28], and it uses a stepwise denoising method to decode the atom coordinates of the whole complex structure with a diffusion transformer. However, most researchers in academia and industry cannot afford such computational costs. Note that AlphaFold3 was not included in the comparison studies in this work because the source code is not available.
IV-C Confidence Assessment
@@
Similarly, Dockformer allocates confidence assessment indicators for each generated complex conformation following the protocol of AlphaFold2. The target confidence of the predicted conformation is estimated via the distance difference test between the predicted and ground truth distances and can be used to describe the binding strengths between proteins and ligands for virtual screening, which is described in Section III-F. To verify the effectiveness of the confidence measures, a scatter diagram of the generated complex conformations with the corresponding confidence indicators and RMSD values is illustrated in Fig. 7(a). We find an apparent linear relationship between confidence and RMSD, represented by the orange line, through a simple linear regression method. The results suggest that higher confidence indicates lower RMSDs of the generated conformations.
@@
In addition, the effectiveness of the confidence indicators was verified by distinguishing the strong and weak binders. The predicted binding poses are allocated positive or negative labels, depending on whether their RMSDs are above or below the threshold of . The confidence indicators were subsequently used to classify these conformations, and the receiver operating characteristic curves are presented in Fig. 7(b). Areas under the curve (AUCs) of 0.7506 and 0.7438 are achieved on two benchmark datasets. These values are much larger than 0.5000, indicating the powerful classification performance of the confidence indicators. Therefore, on the basis of these confidence indicators, Dockformer can be applied to large-scale virtual screening tasks without the need for additional scoring functions.
IV-D Large-scale Virtual Screening Task
@@
In this section, Dockformer is applied to a real-world virtual screening scenario to verify its screening power. Considering that COVID-19 has challenged economic and health care systems worldwide, Dockformer was utilized to screen potential drug candidates for this disease with high transmission and mortality rates. The main protease , whose binding site is highly conserved among all coronaviruses, was selected as the target protein and can serve as a drug target for the design of broad-spectrum inhibitors [37]. Previous studies revealed that the Michael acceptor inhibitor can specifically inhibit the of multiple coronaviruses, including SARS-CoV and MERS-CoV, and has shown potent antiviral activity against infectious bronchitis virus in animal models [38, 39]. The binding pose of and N3 is illustrated in Fig. 8(a). Hydrophobic interactions evidently exist between the residues THR25, MET165, HIS41, and GLN189 and the inhibitor N3, and hydrogen bonds are formed between the residues THR25, GLU166, HIS41, and N3. A large-scale bioactivity database named ChEMBL is used for screening [40]. This database contains more than 1.2 million compounds after filtering the molecules whose molecular weights are greater than 400 . Traditional optimization-based docking approaches require more than one year to screen all the compounds, whereas Dockformer completes such screening tasks in less than 48 hours, highlighting its high efficiency.
@@
Table LABEL:vs presents the top 20 compounds with the highest confidence. The compound named CHEMBL1571559 ranks first, and it forms hydrogen bonds with the residue GLU166 and hydrophobic interactions with the residues MET165 and GLN189 in , as shown in Fig. 8(b). CHEMBL1571559 and N3 exhibit the same interaction patterns. Similar observations can also be found for the compound named CHEMBL277134 in Fig. 8(c). This compound forms hydrogen bonds with residue HIS41 and hydrophobic interactions with residues MET165 and GLN189. In addition, a well-characterized serotonin antagonist named cinanserin can inhibit SARS by forming cation– interactions with the benzene rings of the residues HIS41 and GLU166 [41], [42]. The same interactions can also be observed for the compound named CHEMBL277789, which not only forms hydrogen bonds with the residues ARG188, GLU166, CYS145, and SER144 but also has stacking with the ring structure of the residue HIS41, as shown in Fig. 8(d). These results indicate the potential screening power of Dockformer in practical applications.
@@
IV-E Physical Plausibility Correction
@@
The major disadvantage of DL-based docking methods, including Dockformer, is that their predicted complex conformations might lack physical plausibility. These algorithms directly generate the 3D coordinates of each ligand atom instead of the translation, orientation and torsion of the ligands, intentionally increasing the degree of freedom of docking problems but inevitably distorting the inherent topological structures of the molecules. To address this issue, three postprocessing methods, namely, point cloud fitting-based alignment (PCF), force-field optimization (FF) and energy minimization (EM), have been proposed to refine the predicted binding conformations. Specifically, PCF uses a distance geometry-based method to apply the transformations to the generated conformation. FF employs an iterative process to optimize the conformations regardless of the protein pocket structure; thus, it fails to guarantee intermolecular validity, generally leading to overlap and steric hindrance between ligand and protein atoms. EM modifies the ligand poses by minimizing the binding energy of the ligand–protein complex, considering the rigid structure of the protein pocket.
@@
Fig. 9(a) shows a visualization of binding poses modified by different methods. There are apparent topological distortions in the raw predicted conformation compared with the ground truth (GT) structure. FF fails to correct incorrect topological structures. Although PCF provides the correct local structures, it dramatically changes the original torsion of the ligand, influencing the intermolecular interactions between the ligand and the protein pocket. EM can provide a physically plausible conformation and maintain the original translation, orientation and torsions of the ligands as much as possible. Figs. 9(b) and 9(c) present the performance of these algorithms on both the PDBbind core set and PoseBusters dataset.
@@
The binding poses predicted by Dockformer led to bond length errors of 0.1907 and 0.2288 and angle errors of 0.1572 and 0.2004 on the PDBbind core set and PoseBusters dataset, respectively. PCF obtained the lowest errors for both measures among the postprocessing methods. FF produced unreasonable local structures, leading to higher errors than the other strategies did. Regarding protein–ligand distance restrictions and volume overlaps, the predicted binding modes from Dockformer have correct rates of 0.88 and 0.99 on the PDBbind core set and 0.76 and 0.97 on the PoseBusters dataset. However, these two measures decrease significantly after the PCF and FF strategies are applied, especially for distance restrictions. The EM method maintained the highest accuracy on both benchmarks since the rigid structure of the protein pocket is a mandatory requirement to optimize the ligand conformations. In summary, the EM strategy can effectively refine the binding poses predicted by Dockformer to guarantee physical plausibility.
IV-F Ablation Studies
1) Impact of the Structure Module: To prove the effectiveness of the structure module, the performance of Dockformer Raw was evaluated and compared with that of the baseline model. Dockformer Raw uses a gradient descent method to generate the binding pose instead of the structure module, which is based on the predicted distance matrices. As presented in Table II, Dockformer Raw achieves a success rate of 84.91% and an RMSD of on the PDBbind core set. These values are much lower than those of Dockformer. In addition, the gradient descent method updates the ligand coordinates iteratively, leading to more time consumption during optimization. As shown in Fig. 6, Dockformer Raw requires 11.32 seconds per ligand, which implies much greater computational complexity than Dockformer does. The predicted distance distributions between ligand and protein atoms derived from the pairwise representations of Dockformer and Dockformer Raw are compared in Fig. 10(a). Compared with the ground-truth distance map, Dockformer Raw fails to characterize specific interactions, such as the relationship between ligand atoms 3 and 4 and receptor atoms 36 to 78, as well as the relationships between ligand atom 10 and receptor atoms 54 to 78. These results suggest that the structure module can capture more potential interactive knowledge to generate accurate binding poses than can regular optimization algorithms.
@@ IV-F Ablation Studies
2) Impact of Multimodality Information: Compared with the previous transformer-based docking approaches, Dockformer uses multimodal molecular information to enhance the model capabilities via the following strategies. First, as the conventional self-attention mechanism lacks positional information, the positional embeddings regarding the 1D sequence information are added to each token in the nature language process field. Similarly, learnable global position embeddings, which reflect the spatial information of 3D atomic coordinates, are added to atom embeddings for both ligands and proteins in Dockformer. As expected, the success rate and RMSD of the main architecture without global position embeddings remarkably decrease compared with those of the baseline model (76.49% 90.53% and 1.78 1.14, respectively), implying the effectiveness of global position embeddings. Second, molecular structures contain inherent topologies or local rigid fragments, which can be regarded as both 2D graphs and 3D point clouds. To fully use the multimodal information, Dockformer concatenates the 2D graph information of connection and bond features and the 3D geometric information of interatomic distance features to generate pair representations. As shown in Table II, without using multimodal information, the docking success rate and RMSD of the Dockformer decrease to 84.56% and , respectively. In addition, the multimodal representations are then incorporated to update the atom embeddings as the bias of the talking-head attention mechanism. Removing the talking-head attention changes the success rate and RMSD to 85.96% and respectively, emphasizing the great importance of multimodality information in Dockformer.

### v4

@@ IV-B Prediction Performance on Benchmark Datasets
In this section, the docking accuracy of Dockformer is evaluated on the PDBbind core set and PoseBusters dataset and compared with that of five commonly used optimization-based docking algorithms, namely, GOLD [6], Glide [32], LeDock [33], AutoDock Vina [34] and Mini Vina, and nine state-of-the-art DL-based docking approaches, i.e., DeepDock [20], EquiBind [13], TankBind [21], DiffDock [15], Uni-Mol [14], KarmaDock [22], GAABind [23], CarsiDock [24] and Umol [35].
@@
DL-based docking methods are susceptible to the input training sample distribution, and changing the search space dramatically decreases the model capacity. Therefore, the trained models provided by the official repositories with the default search spaces are used in the comparison study. Specifically, since EquiBind, DiffDock, TankBind and Umol are trained for blind docking tasks, their search spaces cover the entire crystal protein. The binding pockets of Uni-Mol, GAABind and Dockformer are defined as the protein residues within the range of from any heavy atom of a crystal ligand, whereas those of KarmaDock and CarsiDock are considered the protein residues within the range of and , respectively. DeepDock considers the protein surface mesh nodes within of any crystal ligand atom as inputs. In addition, the number of binding pocket boxes is set to for all conventional docking methods. The root mean square deviation (RMSD), which measures the geometric similarity between the predicted binding conformations and the crystal structures of the ligands, is used to evaluate the docking approaches. Generally, binding pose predictions are considered successful when their RMSDs are below the threshold of [36].
@@
As illustrated in Fig. 5(a), Dockformer achieves the highest docking success rates of 90.53% and 82.71% on the PDBbind core set and PoseBusters dataset, respectively. These rates are higher than those of all the baselines. CarsiDock is the second-best model, with slight success rate decreases of 1.76% and 7.01% on the two benchmarks, respectively. TankBind, EquiBind and DiffDock perform relatively poor because they are originally trained to solve blind docking tasks. In addition, the docking performance on the PoseBusters dataset is worse for all approaches but especially for DL-based methods. For example, KarmaDock achieves a success rate of 83.86% using the PDBbind core set and 46.73% using PoseBusters. This finding implies that some DL-based methods may not generalize well to unseen data because PoseBusters contains only complexes released since 2021. However, the accuracy of Dockformer ranks first on this dataset, suggesting its strong generalizability. Interestingly, most optimization-based docking methods achieve satisfactory and robust performance on both datasets. These methods are superior to those of most DL-based methods. LeDock achieves the second-best accuracy and is slightly inferior to Dockformer using PoseBusters. As a more comprehensive and effective method, the cumulative frequencies of the binding poses with the RMSDs from their corresponding crystal ligands for all the docking methods are plotted in Fig. 5(b). We find that the success rates of binding poses generated by GOLD and LeDock are higher than those of most DL-based docking approaches. However, Dockformer still performs very competitively with different RMSD cutoffs on both benchmarks, indicating its obvious superiority in terms of accuracy.
IV-C Computational Complexity versus Accuracy
@@
In addition to accuracy, computational complexity is an important performance indicator that requires attention, especially when the screening compound library becomes extremely large. Most docking methods cannot traverse the entire library within an acceptable running time. As depicted in Fig. 6, Dockformer yields the highest docking accuracy and requires the least amount of time using the PoseBusters dataset among all the docking methods. Although Dockformer has the largest DL network, end-to-end binding conformation generation results in low computational complexity in the inference process. Both LeDock and CarsiDock can also achieve competitive docking accuracies but with a long inference time. Thus, these approaches are unsuitable for LSVS tasks. Moreover, most DL-based docking approaches are more efficient than optimization-based approaches, but they sacrifice accuracy. In addition, the docking accuracy of DL-based methods improves as the size of the network architecture increases.
@@
The computational cost of optimization-based docking methods is substantially greater than that of DL-based methods. The adopted stochastic optimization algorithms require more computing resources to determine the position, orientation, and torsion angles of each ligand conformation according to the scoring functions. The DL-based methods, including EquiBind, TankBind, Uni-Mol, GAABind and CarsiDock, use DL algorithms to construct intra- and intermolecular distance maps and then adopt optimization algorithms to calculate the coordinates of ligand atoms for binding pose generation. These methods cannot abandon independent optimization procedures for each ligand, which are faster than those of optimization-based docking methods but still time-consuming because of the iterative gradient descent processes. Only Dockformer and KarmaDock use end-to-end network architecture modules to generate binding conformations in a batch fashion. These modules have prominent advantages in docking efficiency. A modified version of Dockformer termed Dockformer Raw is also evaluated in this experiment. Dockformer Raw uses the same geometry optimization strategy as that used in TankBind to predict binding conformations.
@@
The experimental results demonstrate that Dockformer Raw performs worse in terms of both accuracy and inference time. This finding highlights that the superiority of end-to-end structures can effectively mitigate the computational burden of iterative optimization. Notably, AlphaFold3 achieves a success rate of 90.2% on the PoseBusters dataset, a rate much higher than that achieved by any of the docking methods mentioned in our experiments. However, even in the case of the fewest number of tokens, the inference time of AlphaFold3 is 22 seconds on 16 A100 graphics processing units (GPUs). This time is much longer than that of most competitors because the model framework of AlphaFold3 is much larger [28], and it uses a stepwise denoising method to decode the atom coordinates of the whole complex structure with a diffusion transformer. However, most researchers in academia and industry cannot afford such computational costs.
IV-D Confidence Assessment
@@
Similarly, Dockformer allocates confidence assessment indicators for each generated complex conformation following the protocol of AlphaFold2. The target confidence of the predicted conformation is estimated via the distance difference test between the predicted and ground-truth distances and can be used to describe the binding strengths between proteins and ligands for virtual screening, which is described in Section III-F. To verify the effectiveness of the confidence measures, a scatter diagram of the generated complex conformations with the corresponding confidence indicators and RMSD values is illustrated in Fig. 7(a). We find an apparent linear relationship between confidence and RMSD, represented by the orange line, through a simple linear regression method. The results suggest that higher confidence indicates lower RMSDs of the generated conformations.
@@
In addition, the effectiveness of the confidence indicators is verified by distinguishing the strong and weak binders. The predicted binding poses are allocated positive or negative labels, depending on whether their RMSDs are above or below the threshold of . The confidence indicators are subsequently used to classify these conformations, and the receiver operating characteristic curves are presented in Fig. 7(b). Areas under the curve (AUCs) of 0.7506 and 0.7438 are achieved on two benchmark datasets. These values are much larger than 0.5000, indicating the powerful classification performance of the confidence indicators. Therefore, on the basis of these confidence indicators, Dockformer can be applied to large-scale virtual screening tasks without additional scoring functions.
IV-E Large-scale Virtual Screening Task
@@
In this section, Dockformer is applied to a real-world virtual screening scenario to verify its screening power. Considering that COVID-19 has challenged economic and healthcare systems worldwide, Dockformer is utilized to screen potential drug candidates for this disease with high transmission and mortality rates. The main protease , whose binding site is highly conserved among all coronaviruses, is selected as the target protein and can serve as a drug target for the design of broad-spectrum inhibitors [37]. Previous studies revealed that the Michael acceptor inhibitor can specifically inhibit the of multiple coronaviruses, including SARS-CoV and MERS-CoV, and has shown potent antiviral activity against infectious bronchitis virus in animal models [38, 39]. The binding pose of and N3 is illustrated in Fig. 8(a). Hydrophobic interactions evidently exist between the residues THR25, MET165, HIS41, and GLN189 and the inhibitor N3, and hydrogen bonds are formed between the residues THR25, GLU166, HIS41, and N3. A large-scale bioactivity database named ChEMBL is used for screening [40]. This database contains more than 1.2 million compounds after filtering the molecules whose molecular weights are greater than 400 . Traditional optimization-based docking approaches require more than one year to screen all the compounds, whereas Dockformer completes such screening tasks in less than 48 hours, highlighting its high efficiency.
@@
Table LABEL:vs presents the top 20 compounds with the highest confidence. The compound named CHEMBL1571559 ranks first, and it forms hydrogen bonds with the residue GLU166 and hydrophobic interactions with the residues MET165 and GLN189 in , as shown in Fig. 8(b). CHEMBL1571559 and N3 exhibit the same interaction patterns. Similar observations can also be found for the compound named CHEMBL277134 in Fig. 8(c). This compound forms hydrogen bonds with residue HIS41 and hydrophobic interactions with residues MET165 and GLN189. In addition, a well-characterized serotonin antagonist named cinanserin can inhibit SARS by forming cation– interactions with the benzene rings of the residues HIS41 and GLU166 [41], [42]. The same interactions can also be observed for the compound named CHEMBL277789, which not only forms hydrogen bonds with the residues ARG188, GLU166, CYS145, and SER144 but also has stacking with the ring structure of the residue HIS41, as shown in Fig. 8(d).
To sum up, these results validate the potential screening power of Dockformer in practical applications, considering it can efficiently screen a large-scale molecular library and precisely identify candidate compounds similar to the known inhibitor N3. That showcases its ability to find potential drug candidates within a vast chemical space quickly and accurately, underscoring its potential utility in accelerating drug discovery efforts.
@@
IV-F Physical Plausibility Correction
@@
The major disadvantage of DL-based docking methods, including Dockformer, is that their predicted complex conformations might lack physical plausibility. These algorithms directly generate the 3D coordinates of each ligand atom instead of the translation, orientation and torsion of the ligands, intentionally increasing the degree of freedom of docking problems but inevitably distorting the inherent topological structures of the molecules. To address this issue, three postprocessing methods—point cloud fitting-based alignment (PCF), force-field optimization (FF), and energy minimization (EM)—are proposed to refine the predicted binding conformations. Specifically, PCF uses a distance geometry-based method to apply the transformations to the generated conformation. FF employs an iterative process to optimize the conformations regardless of the protein pocket structure. Thus, it fails to guarantee intermolecular validity, generally leading to overlap and steric hindrance between ligand and protein atoms. EM modifies the ligand poses by minimizing the binding energy of the ligand–protein complex, considering the rigid structure of the protein pocket.
@@
Fig. 9(a) shows a visualization of binding poses modified by different methods. There are apparent topological distortions in the raw predicted conformation compared with the ground-truth structure. FF fails to correct incorrect topological structures. Although PCF provides the correct local structures, it dramatically changes the original torsion of the ligand, influencing the intermolecular interactions between the ligand and the protein pocket. EM can provide a physically plausible conformation and maintain the original translation, orientation and torsions of the ligands as much as possible. Figs. 9(b) and 9(c) present the performance of these algorithms on both the PDBbind core set and PoseBusters dataset.
@@
The binding poses predicted by Dockformer lead to bond length errors of 0.1907 and 0.2288 and angle errors of 0.1572 and 0.2004 on the PDBbind core set and PoseBusters dataset, respectively. PCF obtaines the lowest errors for both measures among the postprocessing methods. FF produces unreasonable local structures, leading to higher errors than the other strategies did. Regarding protein–ligand distance restrictions and volume overlaps, the predicted binding modes from Dockformer achieve correct rates of 0.88 and 0.99 on the PDBbind core set, and 0.76 and 0.97 on the PoseBusters dataset. However, these two measures decrease significantly after the PCF and FF strategies are applied, especially for distance restrictions. The EM method maintains the highest accuracy on both benchmarks since the rigid structure of the protein pocket is a mandatory requirement to optimize the ligand conformations. In summary, the EM strategy can effectively refine the binding poses predicted by Dockformer to guarantee physical plausibility.
@@
IV-G Ablation Studies
1) Impact of the Structure Module: To prove the effectiveness of the structure module, the performance of Dockformer Raw is evaluated and compared with that of the baseline model. Dockformer Raw uses a gradient descent method to generate the binding pose instead of the structure module, which is based on the predicted distance matrices. As presented in Table II, Dockformer Raw achieves a success rate of 84.91% and an RMSD of on the PDBbind core set. These values are much lower than those of Dockformer. In addition, the gradient descent method updates the ligand coordinates iteratively, leading to more time consumption during optimization. As shown in Fig. 6, Dockformer Raw requires 11.32 seconds per ligand, which implies much greater computational complexity than Dockformer does. The predicted distance distributions between ligand and protein atoms derived from the pairwise representations of Dockformer and Dockformer Raw are compared in Fig. 10(a). Compared with the ground-truth distance map, Dockformer Raw fails to characterize specific interactions, such as those between ligand atoms 3-4 and receptor atoms 36-78, as well as the interaction between ligand atom 10 and receptor atoms 54-78. These results suggest that the structure module can capture more potential interactive knowledge to generate accurate binding poses than regular optimization algorithms can.
2) Impact of Multimodality Information: Compared with the previous transformer-based docking approaches, Dockformer uses multimodal molecular information to enhance the model capabilities via the following strategies. First, as the conventional self-attention mechanism lacks positional information, the positional embeddings for the 1D sequence information are added to each token in the natural language process field. Similarly, learnable global position embeddings, which reflect the spatial information of 3D atomic coordinates, are added to atom embeddings for both ligands and proteins in Dockformer. As expected, the success rate and RMSD of the main architecture without global position embeddings remarkably decrease compared with those of the baseline model (76.49% 90.53% and 1.78 1.14, respectively), implying the effectiveness of global position embeddings. Second, molecular structures contain inherent topological structures and local rigid fragments, which can be interpreted as both 2D graphs and 3D point clouds. To fully use the multimodal information, Dockformer concatenates the 2D graph information of connection and bond features and the 3D geometric information of interatomic distance features to generate pair representations. As shown in Table II, without using multimodal information, the docking success rate and RMSD of the Dockformer decrease to 84.56% and , respectively. In addition, the multimodal representations are then incorporated to update the atom embeddings as the bias of the talking-head attention mechanism. Removing the talking-head attention changes the success rate and RMSD to 85.96% and respectively, emphasizing the great importance of multimodality information in Dockformer.